# Bootstrap Dynamic-Aware 3D Visual Representation for Scalable Robot Learning

### v2

@@ 1 Introduction
Visual pre-training has significantly advanced robotic manipulation by decoupling representation learning from policy learning: Models acquire general visual priors from large-scale datasets before transferring them to downstream control tasks [RPR, hrp, sensorimotor, real-pre, unbiased-pre-training, exploring, learningact]. This paradigm enables richer feature learning and facilitates cross-domain transfer. Recent 2D approaches (e.g., MVP [mvp_ref], R3M [r3m_ref], VIP [vip_ref]) leverage large-scale image or video corpora for training and have substantially improved manipulation performance.
@@
However, while 3D point clouds inputs offer more geometric information for better efficiency and generalization [rvt, whole, points], 3D visual pre-training for robotic manipulation remains underexplored; meanwhile, standardized pipelines and generative tools are rapidly expanding 3D data quality and scale. [robomind, rh20t]. These trends motivate a central question: Can we design a scalable 3D visual pre-training framework specially tailored to robotic manipulation?
@@
Point clouds offer a natural format for 3D data, yet their irregular structure, sparsity, and permutation invariance complicate spatial reasoning and semantic abstraction [pointnet_ref]. Although many 3D pre-training methods have effectively overcome these challenges and demonstrate strong performance on tasks such as object recognition [pointbert, strl] and segmentation [pointcontrast], they still exhibit suboptimal transfer to robotic manipulation, and even lagging behind 2D foundation models in manipulation benchmarks [lift3d]. We identify that current 3D pre-training approaches overlook two critical aspects:
@@
Lack of dynamics awareness. Robotic manipulation is inherently sequential, where each action brings about deliberate changes in the state. However, most existing 3D pre-training frameworks use supervision signals derived solely from the same frame as their inputs. This design overlooks the temporal continuity and causal dependencies underlying state transitions, resulting in a representation space that lacks coherent temporal structure and fails to capture dynamic or causal relationships across states.
@@
Lack of manipulation-relevant abstraction. While many 3D pre-training approaches focus on holistic scene reconstruction, they often capture background details irrelevant to control [univla]. However, incorporating such dense representations can mislead policy networks, as they may divert attention from task-critical elements. In contrast, effective manipulation necessitates abstractions that prioritize actionable object regions and interaction dynamics, rather than striving for full-scene fidelity.
@@
Robotic manipulation unfolds as state–action–state trajectories rich in supervisory signal.Yet scalable, task-general visual representation learning often omits these labels, creating a tension: how can a model internalize dynamics without explicit action or transition annotations? We posit that dynamics can emerge through latent-space objectives.Recent advances in latent-predictive learning (e.g., V-JEPA 2 [v-jepa2]) show that aligning high-level features without pixel-level reconstruction yields broadly generalizable representations.Building on this insight, we introduce AFRO, a self-supervised 3D pre-training framework that learns dynamic-aware, manipulation-relevant representations from point clouds. AFRO aims to (i) learn dynamic-aware representations to model the spatiotemporal relationships across states, and (ii) acquire manipulation-relevant abstractions without explicit reconstruction.
@@
Concretely, AFRO integrates inverse and forward dynamics models (IDM/FDM) to encode state–action–state transitions directly in a latent space. As shown in Figure LABEL:fig1:teaser, consecutive point clouds are encoded into feature vectors; the IDM infers a latent action from current and subsequent features, while the FDM predicts the next features given the current features and the inferred latent action. State prediction is performed entirely in the feature space, where the online encoder is trained using a VICReg [vicreg] loss against an exponential moving average(ema) target encoder. To handle multimodal future outcomes, we formulate feature prediction as a diffusion-based generation process that encourages upstream modules to supply informative conditions. Through this dynamics-driven learning, the feature encoder is guided to attend to regions with motion changes, thereby encoding representations that are more relevant to manipulation.
@@
To prevent latent action shortcutting and feature leakage in IDM learning, we input feature differences instead of raw feature pairs, forcing the model to reason over action-driven change rather than memorize states. We further introduce inverse-consistency supervision: the same pipeline is applied in reverse to explain previous states from future observations. By requiring the latent action to account for both forward prediction and backward explanation, we curb degenerate solutions and stabilize representation learning.
@@
We evaluate AFRO on 16 simulated and 4 real-world manipulation tasks spanning diverse skills (e.g., sliding, pressing, pushing). Across benchmarks, AFRO consistently outperforms strong 2D/3D pretraining baselines and imitation-from-scratch, with advantages that scale with both data volume and domain diversity. Pretraining on the large-scale RH20T [rh20t] dataset further strengthens downstream results, yielding robust gains across all tasks.
@@
Our core contributions are: (1) We propose a 3D visual pretraining framework for robotic manipulation that learns dynamics-aware representations directly in latent space, modeling future state uncertainties with a generative diffusion process while avoiding explicit reconstruction. (2) To our knowledge, we are the first to introduce latent actions into 3D visual learning, and design feature-differencing and inverse-consistency supervision to prevent shortcut learning and feature leakage, thereby enhancing representation quality and stability. (3) Extensive simulation and real-robot experiments show that our method outperforms a range of advanced competing baselines; visualizations and ablations further validate the contribution of each component.
@@ 2.1 Visual Pretraining for Robot Manipulation
Visual pretraining significantly enhances data efficiency in robotic manipulation by learning transferable representations from large-scale visual data [embodiedmae, ar_vrm]. Early 2D approaches such as R3M [r3m_ref], MVP [mvp_ref], and VIP [vip_ref] leverage contrastive or masked modeling techniques to extract semantic pixel-level features from images. However, these methods are confined to image-space reasoning and lack explicit 3D geometric awareness, which is a crucial aspect for accurately perceiving spatial in real-world environments.
Recent work has thus shifted toward 3D visual pretraining to better capture scene geometry. For instance, 3D-MVP [3d-mvp] extends MVP by incorporating multi-view masked autoencoding on large-scale 3D data; SUGAR [sugar] jointly learns semantic and affordance representations from point clouds, thereby improving zero-shot manipulation capabilities. However, most existing 3D methods still overlook the inherent temporal dynamics that characterize robotic manipulation. Although FVP [fvp_ref] advances this direction with diffusion-based future prediction, it remains dependent on explicit scene reconstruction. In contrast, our approach AFRO introduces action-free 3D pretraining that learns dynamic-aware representations directly in latent space, effectively bridging 3D perception and embodied behavior without relying on action labels or reconstruction.
@@ 2.2 Latent Actions for Robot Learning
Latent actions encode state transitions into low-dimensional variables, providing an alternative to expensive action annotations. [latent-supervision, taco, latent-behavior, adaworld, allshire2021laser] Early methods used inverse dynamics models (IDM) for pretraining, while later approaches jointly learned IDM and forward dynamics models (FDM) in visual spaces to improve sample efficiency. [villa, AMPLIFY, latent-world, jif] Current research falls into three categories: (1) Using latent actions as policy supervision from unlabeled videos, as in CLAM [clam] and CoMo [como]; (2) Vision-language-action pretraining—e.g., LAPA [lapa] discretizes frame changes into tokens for efficient regression, and UniVLA [univla] learns task-centric latent actions in DINO space for cross-embodiment generalization; (3) Reinforcement learning with constrained action spaces to address OOD issues, as in PLAS [plas].
Recently, Dynamo [dynamo_ref] leverages latent actions in visual pretraining to capture dynamic changes in manipulation, enabling it to achieve strong performance transfer on downstream tasks. However, these methods remain confined to 2D visual spaces. We adapt latent actions to 3D pretraining and mitigate IDM’s tendency to overfit future features by: (i) using feature differencing rather than frame stacking as IDM input, and (ii) introducing inverse-consistency supervision, whereby inverse latent actions are constructed to enable the FDM to predict historical states from future ones, thereby improving reversibility and temporal consistency.
Figure 1: Overall framework of our method. (a) Predict Future: Given point clouds $\mathcal{P}_{t}$ and $\mathcal{P}_{t+k}$, the online encoder $f_{\phi}$ encodes both to obtain $\mathbf{z}_{t}$ and $\mathbf{z}_{t+k}$. The inverse dynamics model $g_{\psi}$ takes the difference $(\mathbf{z}_{t+k}-\mathbf{z}_{t})$ to infer the forward latent action $\boldsymbol{\alpha}_{t\rightarrow t+k}$. The target encoder $f_{\xi}$ (EMA-updated) encodes only $\mathcal{P}_{t+k}$ to yield the teacher target $\tilde{\mathbf{z}}_{t+k}$. The forward dynamics model $h_{\theta}$ predicts the future feature by mapping $(\mathbf{z}_{t},\boldsymbol{\alpha}_{t\rightarrow t+k})\mapsto\hat{\mathbf{z}}_{t+k}$ and aligning it to $\tilde{\mathbf{z}}_{t+k}$, which explicitly drives $f_{\phi}$ to learn dynamics-aware representations. (b) Predict History: Different from (a), using the difference $(\mathbf{z}_{t}-\mathbf{z}_{t+k})$, $g_{\psi}$ infers the backward latent action $\boldsymbol{\alpha}_{t+k\rightarrow t}$; $f_{\xi}$ encodes only $\mathcal{P}_{t}$ to obtain $\tilde{\mathbf{z}}_{t}$; then $h_{\theta}$ maps $(\mathbf{z}_{t+k},\boldsymbol{\alpha}_{t+k\rightarrow t})\mapsto\hat{\mathbf{z}}_{t}$ and aligns it with $\tilde{\mathbf{z}}_{t}$. Other steps are symmetric to (a) and omitted for brevity. (c) Notation Summary: $\mathcal{P}_{t}$ point cloud; $f_{\phi}$ online encoder; $f_{\xi}$ target encoder; $\mathbf{z}$ student feature; $\tilde{\mathbf{z}}$ teacher feature; $g_{\psi}$ inverse dynamics model; $\boldsymbol{\alpha}$ latent action; $h_{\theta}$ forward dynamics model.
@@ 3 Preliminary
We aim to learn dynamics-aware 3D representations from unlabeled point cloud demonstrations. Our approach builds on a latent dynamics modeling paradigm, consisting of an online visual encoder $f_{\phi}$, an inverse dynamics model (IDM) $g_{\psi}$, and a forward dynamics model (FDM) $h_{\theta}$. Given consecutive point clouds $\mathcal{P}_{t}$ and $\mathcal{P}_{t+k}$, the encoder first extracts their features $\mathbf{z}_{t}$ and $\mathbf{z}_{t+k}$. The IDM then infers a latent action that explains the transition:
@@
which is intended to capture the underlying motion. The FDM uses this action to predict the future feature:
@@
enabling state prediction entirely in the latent space. An EMA-updated target encoder $f_{\xi}$ provides stable regression targets $\tilde{\mathbf{z}}$, avoiding explicit reconstruction.
While this framework offers a principled starting point, we identify two critical limitations in its naive implementation. First, the IDM is prone to feature leakage instead of learning meaningful transitions, as it may exploit shortcuts by directly encoding information from $\mathbf{z}_{t+k}$, leading to degenerate latent actions that fail to represent dynamics. Second, standard FDMs are typically deterministic and struggle to capture the multimodal uncertainty of real-world interactions, often predicting averaged, unrealistic futures.
To overcome these issues, we introduce two core innovations in Sec. 4:
@@
    A feature-differencing and reverse consistency framework for latent action learning, which prevents feature leakage and enhances temporal coherence (Sec. 4.1).
@@
    A diffusion-based forward dynamics model that captures multimodal future distributions, enabling robust and diverse state prediction (Sec. 4.2).
@@ 4 Method
Figure 2: Given the differential feature $\big(\mathbf{z}_{t+k}-\mathbf{z}_{t}\big)$ or $\big(\mathbf{z}_{t}-\mathbf{z}_{t+k}\big)$, the inverse dynamics model $g_{\psi}$ passes it through two stacked Linear–GELU layers followed by a final Linear projection to produce the latent action $\boldsymbol{\alpha}_{t\rightarrow t+k}$ or $\boldsymbol{\alpha}_{t+k\rightarrow t}$, which encodes the implicit motion between consecutive states.
@@ 4.1 Latent Action Modeling
We therefore model changes in representation rather than the representations themselves. Concretely, instead of feeding both $\mathbf{z}_{t}$ and $\mathbf{z}_{t+k}$ into the inverse model, we take their difference and learn latent actions directly from the displacement:
@@
which aligns the two directions conceptually and constrains latent actions to be causally consistent. Taken together, differencing plus inverse-consistency provides a lightweight but effective recipe for learning latent actions without explicit labels.
@@ 4.2 Forward Dynamics with Diffusion Transformer
The future is inherently uncertain and often multimodal, as occlusions and stochastic interactions give rise to many plausible outcomes. Conventional MLP- or transformer-based predictors struggle to capture such uncertainty and tend to average over multiple modes. Instead, we formulate forward prediction as a conditional denoising process: given the current feature $\mathbf{z}_{t}$, the latent action $\boldsymbol{\alpha}_{t\rightarrow t+k}$, and the diffusion step $\tau$, the forward dynamics model $h_{\theta}$—implemented as a diffusion transformer [dit] (DiT) with AdaLN-Zero conditioning—maps a noisy sample along with these conditioning signals to a clean future latent representation.
@@
Direct denoising (student). At training and inference, we take the noisy future feature and produce the clean target in a single pass:
$\hat{\mathbf{z}}_{t+k}=h_{\theta}\!\big(\mathbf{z}^{(\tau)}_{t+k},\;\mathbf{z}_{t},\;\boldsymbol{\alpha}_{t\rightarrow t+k},\;\tau\big),$ (7)
so the noisy input and condition $(\mathbf{z}_{t},\boldsymbol{\alpha}_{t\rightarrow t+k},\tau)$ flow through DiT blocks and a projector to yield $\hat{\mathbf{z}}_{t+k}$ in one step.
@@
4.3 Variance-Regularized Representation Matching
Self-supervised latent learning can easily collapse into trivial constant embeddings. To prevent this, we adopt a Variance–Invariance–Covariance Regularization (VICReg) [vicreg] loss that preserves the statistical diversity. Given predicted features $\hat{\mathbf{z}}$ and teacher target features $\tilde{\mathbf{z}}$ from the EMA encoder $f_{\xi}$, the loss combines three complementary terms:

### v3

@@ 1 Introduction
Visual pre-training has significantly advanced robotic manipulation by decoupling representation learning from policy learning: models first acquire visual priors from large-scale datasets and then transfer them to downstream control tasks [19, 40, 36, 37, 9, 20, 56]. This paradigm enables richer features and cross-domain transfer. Recent 2D approaches (e.g., MVP [49], R3M [27], VIP [26]) leverage large-scale image or video corpora and substantially improve manipulation performance.
@@
However, although 3D point cloud inputs provide richer geometric information and stronger generalization [14, 25, 66], 3D visual pre-training for robotic manipulation remains underexplored, even as standardized pipelines and generative tools rapidly expand the scale and quality of 3D data [46, 11]. These trends motivate a central question: Can we design a scalable 3D visual pre-training framework tailored to robotic manipulation?
@@
Point clouds provide a natural representation of 3D scenes, yet their irregular structure, sparsity, and permutation invariance complicate spatial reasoning and semantic abstraction [33]. Although many 3D pre-training methods perform well on object recognition [58, 17] and segmentation [50], they transfer poorly to robotic manipulation and can even lag behind 2D foundation models on manipulation benchmarks [18]. We identify two key limitations in existing 3D pre-training approaches:
@@
Lack of dynamics awareness. Robotic manipulation is inherently sequential, where each action changes the system state. However, most 3D pre-training frameworks rely on supervision from single frames, ignoring temporal continuity and causal dependencies between states. As a result, the learned representations lack coherent temporal structure and fail to capture dynamic relationships across states.
@@
Lack of manipulation-relevant abstraction. Many 3D methods focus on holistic scene reconstruction and often capture background details irrelevant to control [4]. Such dense representations may mislead policy networks by diverting attention from task-critical elements. In contrast, effective manipulation requires abstractions that emphasize actionable object regions and interaction dynamics.
@@
Robotic manipulation forms state–action–state trajectories rich in supervisory signals. Yet scalable visual representation learning often omits these labels, raising a key question: how can models capture dynamics without explicit action or transition annotations? We argue that dynamics can emerge through latent-space objectives. Recent latent-predictive learning approaches (e.g., V-JEPA 2 [2]) show that aligning high-level features without pixel-level reconstruction can produce generalizable representations.
Building on this, we introduce AFRO, a self-supervised 3D pre-training framework for learning dynamics-aware representations from point clouds in a world-modeling sense [55]. AFRO aims (i) to model spatiotemporal relationships across states and (ii) to learn manipulation-relevant abstractions without explicit reconstruction.
@@
Concretely, AFRO integrates inverse and forward dynamics models (IDM/FDM) to encode state–action–state transitions in latent space. As shown in Figure LABEL:fig1:teaser, consecutive point clouds are encoded into features; the IDM infers a latent action from adjacent states, while the FDM predicts future features conditioned on the current feature and inferred action. State prediction occurs entirely in feature space, where the encoder is trained using a VICReg [3] loss against an exponential moving average (EMA) target encoder. To capture multimodal futures, we model feature prediction as a diffusion-based generation process.
@@
To prevent latent action shortcutting and feature leakage in IDM learning, we input feature differences instead of raw feature pairs, forcing the model to reason about action-driven change rather than memorizing states. We further introduce inverse-consistency supervision, applying the same pipeline in reverse to explain previous states from future observations. By requiring latent actions to support both forward prediction and backward explanation, we avoid degenerate solutions and stabilize representation learning.
@@
We evaluate AFRO on 16 simulated and 4 real-world manipulation tasks spanning skills (e.g., sliding, pressing, pushing). Across benchmarks, AFRO outperforms 2D/3D pretraining baselines and imitation-from-scratch, with gains that scale with both data volume and domain diversity. Pretraining on the large-scale RH20T [11] dataset further improves downstream results across all tasks.
@@
Our core contributions are: (1) We propose a 3D visual pretraining framework for robotic manipulation that learns dynamics-aware representations directly in latent space, modeling future state uncertainty with diffusion while avoiding explicit reconstruction. (2) To our knowledge, we are the first to introduce latent actions into 3D visual learning and design feature-differencing and inverse-consistency supervision to prevent shortcut learning and feature leakage, improving representation quality and stability. (3) Extensive simulation and real-robot experiments demonstrate clear improvements over strong baselines, with visualizations and ablations validating each component.
@@ 2.1 Visual Pretraining for Robot Manipulation
Visual pretraining significantly enhances data efficiency in robotic manipulation by learning transferable representations from large-scale visual data [10, 52]. Early 2D approaches such as R3M [27], MVP [49], and VIP [26] leverage contrastive or masked modeling techniques to extract semantic pixel-level features. However, these methods are largely confined to image-space reasoning and lack explicit 3D geometric awareness, which is crucial for accurate spatial perception in complex real-world environments.
Recent work has shifted toward 3D visual pretraining to better capture scene geometry. For instance, 3D-MVP [34] extends MVP by incorporating multi-view masked autoencoding on large-scale 3D data; SUGAR [5] jointly learns semantic and affordance representations from point clouds, improving zero-shot manipulation capabilities. However, most existing 3D methods still overlook the inherent temporal dynamics of robotic manipulation. Although FVP [15] advances this direction with diffusion-based future prediction, it remains dependent on explicit scene reconstruction. In contrast, our approach AFRO introduces action-free 3D pretraining that learns dynamics-aware representations directly in latent space, bridging 3D perception and embodied behavior without relying on action labels or reconstruction.
Figure 1: Overall framework of our method. (a) Predict Future: Given point clouds $\mathcal{P}_{t}$ and $\mathcal{P}_{t+k}$, the online encoder $f_{\phi}$ produces features $\mathbf{z}_{t}$ and $\mathbf{z}_{t+k}$. The inverse dynamics model $g_{\psi}$ takes the difference $(\mathbf{z}_{t+k}-\mathbf{z}_{t})$ to infer the latent action $\boldsymbol{\alpha}_{t\rightarrow t+k}$. The target encoder $f_{\xi}$ (EMA) encodes $\mathcal{P}_{t+k}$ to obtain $\tilde{\mathbf{z}}_{t+k}$. The forward dynamics model $h_{\theta}$ predicts $(\mathbf{z}_{t},\boldsymbol{\alpha}_{t\rightarrow t+k})\mapsto\hat{\mathbf{z}}_{t+k}$ and aligns it with $\tilde{\mathbf{z}}_{t+k}$, guiding $f_{\phi}$ to learn dynamics-aware representations. (b) Predict History: Using $(\mathbf{z}_{t}-\mathbf{z}_{t+k})$, $g_{\psi}$ infers $\boldsymbol{\alpha}_{t+k\rightarrow t}$. The encoder $f_{\xi}$ encodes $\mathcal{P}_{t}$ to obtain $\tilde{\mathbf{z}}_{t}$, while $h_{\theta}$ predicts $(\mathbf{z}_{t+k},\boldsymbol{\alpha}_{t+k\rightarrow t})\mapsto\hat{\mathbf{z}}_{t}$ aligned with $\tilde{\mathbf{z}}_{t}$. Other steps are symmetric to (a). (c) Notation overview.
@@ 2.2 Latent Actions for Robot Learning
Latent actions encode state transitions into low-dimensional variables, offering an alternative to expensive action annotations [28, 61, 23, 12, 1]. Early work employed inverse dynamics models (IDM) for pretraining, while later approaches jointly learned IDM and forward dynamics models (FDM) in visual spaces to improve sample efficiency [6, 8, 41, 21]. Current research falls into three directions: (1) using latent actions as policy supervision from unlabeled videos, e.g., CLAM [24] and CoMo [53]; (2) vision-language-action pretraining, where LAPA [54] discretizes frame changes into tokens and UniVLA [4] learns task-centric latent actions in DINO space for cross-embodiment generalization; and (3) reinforcement learning with constrained action spaces to address OOD issues, PLAS [65].
Recently, Dynamo [30] incorporates latent actions into visual pretraining to explicitly capture dynamic changes in manipulation. However, existing approaches remain largely limited to 2D visual spaces. In contrast, we introduce latent actions into 3D pretraining and mitigate IDM’s tendency to overfit future features by (i) using feature differencing instead of frame stacking as IDM input, and (ii) introducing inverse-consistency supervision that enables FDM to predict historical states from future observations, improving reversibility and better temporal consistency.
@@ 3 Preliminary
We aim to learn dynamics-aware 3D representations from unlabeled point cloud demonstrations. Our approach adopts a latent dynamics paradigm with an online encoder $f_{\phi}$, inverse dynamics model (IDM) $g_{\psi}$, and forward dynamics model (FDM) $h_{\theta}$. Given consecutive point clouds $\mathcal{P}_{t}$ and $\mathcal{P}_{t+k}$, the encoder extracts features $\mathbf{z}_{t}$ and $\mathbf{z}_{t+k}$. The IDM infers a latent action
@@
and the FDM predicts the future feature
@@
enabling state prediction in latent space, while an EMA-updated encoder $f_{\xi}$ provides stable targets $\tilde{\mathbf{z}}$.
However, two limitations remain. First, the IDM may suffer from feature leakage, where shortcuts encode information from $\mathbf{z}_{t+k}$ and produce degenerate latent actions. Second, standard FDMs are deterministic and struggle with multimodal uncertainty in real-world interactions, often predicting averaged futures. To address these issues, we introduce two components in Sec. 4:
@@
    feature-differencing and reverse-consistency framework for latent action learning that prevents feature leakage and improves temporal coherence (Sec. 4.1).
@@
    diffusion-based forward dynamics model that captures multimodal future distributions (Sec. 4.2).
@@ 4 Method
Figure 2: Given differential features $(\mathbf{z}_{t+k}-\mathbf{z}_{t})$ or $(\mathbf{z}_{t}-\mathbf{z}_{t+k})$, the inverse dynamics model $g_{\psi}$ applies two Linear–GELU layers and a final Linear projection to produce latent actions $\boldsymbol{\alpha}_{t\rightarrow t+k}$ or $\boldsymbol{\alpha}_{t+k\rightarrow t}$, encoding motion across states.
@@ 4.1 Latent Action Modeling
We model changes in representations rather than the representations themselves. Concretely, instead of feeding $\mathbf{z}_{t}$ and $\mathbf{z}_{t+k}$ into the inverse model, we take their difference and learn latent actions directly from the displacement:
@@
which aligns the two directions and constrains latent actions to be causally consistent. Differencing plus inverse-consistency provides a lightweight effective recipe for learning latent actions without explicit labels.
@@ 4.2 Forward Dynamics with Diffusion Transformer
The future is uncertain and multimodal, as occlusions and stochastic interactions produce multiple plausible outcomes. Conventional MLP- or transformer-based predictors struggle to capture such uncertainty and often average over modes. We instead formulate forward prediction as conditional denoising: given the current feature $\mathbf{z}_{t}$, latent action $\boldsymbol{\alpha}_{t\rightarrow t+k}$, and diffusion step $\tau$, the forward dynamics model $h_{\theta}$—implemented as a diffusion transformer [32] (DiT) with AdaLN-Zero conditioning—maps a noisy sample to a clean future latent representation.
@@
Direct denoising (student). The noisy latent is mapped to a clean target in a single pass:
$\hat{\mathbf{z}}_{t+k}=h_{\theta}\!\big(\mathbf{z}^{(\tau)}_{t+k},\;\mathbf{z}_{t},\;\boldsymbol{\alpha}_{t\rightarrow t+k},\;\tau\big).$ (7)
The noisy input and condition $(\mathbf{z}_{t},\boldsymbol{\alpha}_{t\rightarrow t+k},\tau)$ pass through DiT blocks and a projector to produce $\hat{\mathbf{z}}_{t+k}$.
@@
4.3 VICReg Matching
To prevent collapse in self-supervised latent learning, we adopt Variance–Invariance–Covariance Regularization (VICReg) [3]. Given predicted features $\hat{\mathbf{z}}$ and teacher targets $\tilde{\mathbf{z}}$ from the EMA encoder $f_{\xi}$, the loss combines three terms: